Details
-
Type:
Story
-
Status: Done
-
Resolution: Done
-
Fix Version/s: None
-
Component/s: pipe_analysis
-
Labels:None
-
Story Points:4
-
Epic Link:
-
Sprint:DRP F17-5
-
Team:Data Release Production
Description
While several updates have been made to the visit and coadd scripts, the run comparison and color analysis scripts have suffered some bitrot. Please update the visit and coadd run comparison scripts and the coadd color analysis script to current APIs.
Attachments
Issue Links
- mentioned in
-
Page Loading...
Might you have a moment to give this a quick look? I've fixed the butler issue you also came across for the comparison scripts and also made a number of other updates. All scripts seem to be in working order. You can have a look at the current status of the plots on tiger at Princeton at /tigress/HSC/HSC/rerun/lauren/LSST/working/cosmos/plots/. Note that for the "comparison" plots, I'm actually comparing the same run against itself (I didn't have two different recent runs handy, but I will do further testing on Hsin-Fang's weekly processing soon). Of interest
though is for the visitComparison I applied the meas_mosaic results to only one, so these plots in particular show the uberCal correction fields on a per object basis for the mag and astrom comparisons, e.g.:
/tigress/HSC/HSC/rerun/lauren/LSST/working/cosmos/plots/HSC-I/tract-9813/visit-1228/compareVisit-v1228-diff_base_GaussianFlux-sky-gals.png